OpenStudy (anonymous):

2x - y = -2 3y = 6 - 5x Solve for x and y

OpenStudy (anonymous):

2x+2 = y 3y=6-5x 3(2x+2)=6-5x 6x+4=6-5x 6x-5x=6-4 x=2 3y=6-5x 3y=6-5(2) 3y=6-10 3y=-4 3/3y=-4/3 y=4/3 or 1.3333333333333333333333333333333333333333333333